Cardinal Company is considering a project that would require a $2,985,000 investment in equipment with a useful life of five years. At the end of five years, the project would terminate and the equipment would be sold for its salvage value of $400,000. The company’s discount rate is 16%. The project would provide net operating income each year as follows:

  

  Sales $ 2,737,000
  Variable expenses 1,001,000
  Contribution margin 1,736,000
  Fixed expenses:
  Advertising, salaries, and other
    fixed out-of-pocket costs
$ 610,000
  Depreciation 517,000
  Total fixed expenses 1,127,000
  Net operating income $ 609,000

  

Required:

What is the project’s payback period? (Round your answer to 2 decimal places.)

  

  Project’s payback period

Can't get it to come out......help me!

  years
Answer #1

Annual cash flow = 609000+517000 = 1126000

Payback period = Initial investment/Annual cash flow

= 2985000/1126000

Payback period = 2.65 years